# Canonical decompositions of 3-connected graphs

### v2

@@ 1.2. Properties of tri-separations
The implication (2$\to$(1) is clear. For (1$\to$(2) suppose that $G[A]$ contains no cycle, say. We first show that the side $B$ cannot have exactly two vertices. Indeed, if $B$ has size two, then $G[B]$ has maximum degree at most one, and since $(A,B)$ is a tri-separation it follows that $A\cap B$ must be empty. Thus, $A$ and $B$ are the two sides of a cut of size three. Hence some vertex in $B$ has degree at most two in $G$, which contradicts 3-connectivity. So $B$ does not have size two. As we are done otherwise, we from now on assume that the side $B$ contains at least three vertices.
@@
Let any trivial tri-separation be given. By Section 1.2, it is of the form ($(\{v\},V\setminus\{v\})$, say. Then the vertex $v$ has degree three in $G$. Let $(C,D)$ be any strong tri-separation. Since $(C,D)$ is strong, the vertex $v$ is not in $C\cap D$. So it is in precisely one of $C\setminus D$ and $D\setminus C$, say in $C\setminus D$. Then $D\subseteq V\setminus\{v\}$, which gives ($(\{v\},V\setminus\{v\})\leqslant(C,D)$. ∎
@@
Let us show first that no two edges in $S(A,B)$ share an end. For this, let us suppose for a contradiction that two edges $e,f\in S(A,B)$ share an endvertex $v\in A\setminus B$, say. Let $x$ be the remaining element of $S(A,B)$ besides $e,f$ if it is a vertex, and otherwise let $x$ denote the endvertex in $A\setminus B$ of the edge in $S(A,B)$ besides $e,f$. Let $O$ be a cycle in $G[A]$. Since $O$ has at least three vertices, one of them is distinct from $v$ and $x$, and so is not in $B\cup\{v,x\}$. Hence the pair $(A,B\cup\{v,x\})$ is a 2-separation of $G$ with separator equal to $\{v,x\}$. This contradicts the fact that $G$ is 3-connected. ∎
@@
Since $G$ is 3-connected, $v$ has neighbours in $X_{1}\setminus X_{2}$ and in $X_{2}\setminus X_{1}$, and $v$ has degree three. So if (1) fails, we have (2). ∎
@@
Let $(X_{1},X_{2})$ be a mixed 3-separation of a 3-connected graph. We obtain a tri-separation from $(X_{1},X_{2})$ by deleting vertices from $X_{1}$ or $X_{2}$, as follows. For every vertex $v\in X_{1}\cap X_{2}$ that has fewer than two neighbours in some side $X_{i}$, the index $i=:i(v)$ is unique, $v$ has a unique neighbour $x(v)$ in $X_{i}$ that lies in $X_{i}\setminus X_{3-i}$, and $v$ has two neighbours in $X_{3-i}\setminus X_{i}$ by Section 1.2. For both $j\in\{1,2\}$ we obtain $X_{j}^{\prime}$ from $X_{j}$ by deleting all vertices $v$ with $i(v)=j$. Then $(X_{1}^{\prime},X_{2}^{\prime})$ is a tri-separation of $G$. Every vertex $v\in X_{1}\cap X_{2}$ that was removed from some side is not in the separator of $(X_{1}^{\prime},X_{2}^{\prime})$, but instead the edge $\{v,x(v)\}$ is in $S(X_{1}^{\prime},X_{2}^{\prime})$. In this context, we say that $v$ was reduced to the edge $\{v,x(v)\}$. We call $(X_{1}^{\prime},X_{2}^{\prime})$ the reduction of $(X_{1},X_{2})$. Note that the reduction $(X_{1}^{\prime},X_{2}^{\prime})$ is nontrivial if $(X_{1},X_{2})$ is nontrivial.
Let $(A,B)$ be a mixed 3-separation of a graph $G$. A strengthening of $(A,B)$ is a mixed 3-separation $(A^{\prime},B^{\prime})$ that it is obtained from $(A,B)$ by deleting all vertices of $A\cap B$ that have degree three in $G$ and have a neighbour in $A\cap B$ from one of the sides, then taking a reduction.
@@
Every mixed 3-separation $(A,B)$ of a 3-connected graph $G$ has a strengthening $(A^{\prime},B^{\prime})$. Moreover, if there is an edge $uv$ in $G$ with both ends $u,v$ in the separator of $(A,B)$, then we may choose $(A^{\prime},B^{\prime})$ so that $u,v\in B^{\prime}$.
@@
Given $(A,B)$, we obtain $A^{\prime\prime}$ from $A$ by deleting all vertices that lie in $A\cap B$ and have degree three in $G$, and we put $B^{\prime\prime}:=B$. Then we let $(A^{\prime},B^{\prime})$ be the reduction of $(A^{\prime\prime},B^{\prime\prime})$. Suppose now that $uv$ is an edge with ends $u,v\in A\cap B$. Then $u$ and $v$ have two neighbours in $B=B^{\prime\prime}$, so $u,v\in B^{\prime}$. ∎
@@
Clearly 3$\to$4.
[…]
Case 1: the induced subgraph $G[X]$ has no edges. Then $|A\setminus B|\geqslant 2$ and $|B\setminus A|\geqslant 2$. By Section 1.2, there is a strong tri-separation $(A^{\prime},B^{\prime})$ of $(A,B)$ with $A\setminus B\subseteq A^{\prime}$ and $B\setminus A\subseteq B^{\prime}$. Since $A\setminus B$ and $B\setminus A$ have size at least two, so have $A^{\prime}$ and $B^{\prime}$. Thus $(A^{\prime},B^{\prime})$ is nontrivial by Section 1.2.
Case 2: the induced subgraph $G[X]$ contains an edge $x_{1}x_{2}$. Recall that $|G|\geqslant 5$. The only 3-connected graphs on exactly five vertices that have a 3-separator are the 4-wheel and $K_{5}^{-}$ ($K_{5}$ minus one edge). The unique $3$-separator of $K_{5}^{-}$ is a triangle and thus it is the separator of a nontrivial tri-separation. The 4-wheel has two nontrivial strong tri-separations. So it remains to consider the case that $G$ has at least six vertices. By symmetry, we may assume that $|A\setminus B|\geqslant 2$. Since $x_{1}x_{2}$ is an edge of $G$ with both ends in $X=A\cap B$, we find a strong tri-separation $(A^{\prime},B^{\prime})$ of $G$ with $A\setminus B\subseteq A^{\prime}$ and $B\setminus A\subseteq B^{\prime}$ such that $x_{1},x_{2}\in B^{\prime}$. Hence, the side $B^{\prime}$ misses at most one vertex of $B$. As $|B|\geqslant 4$, this gives $|B^{\prime}|\geqslant 3$. We also have $|A^{\prime}|\geqslant|A\setminus B|$, and $|A\setminus B|\geqslant 2$ by assumption. Hence the strong tri-separation $(A^{\prime},B^{\prime})$ is nontrivial by Section 1.2. ∎
1.3. Nested or crossed: analysing corner diagrams
@@
Let $(A,B)$ and $(C,D)$ be mixed-separations of a graph $G$. The following definitions all depend on the context that $(A,B)$ and $(C,D)$ are given. They are supported by Figure 1.2, which is commonly referred to as a ‘corner diagram’.
Figure 1.2. $(A,B)$ and $(C,D)$ cross. Corners are blue, links are red, the centre is yellow.
@@
An edge $e$ in the separator of $(A,B)$ is in the edge-link for $C$ if it is not diagonal and has an endvertex in one of the corners for $C$; that is, in one of the corners for $\{A,C\}$ or $\{B,C\}$. The link for $C$ is the union of the edge-link for $C$ and the vertex set $(A\cap B)\setminus D$. In a slight abuse of notation we will sometimes say things like ‘a vertex of the link $(A\cap B)\setminus D$’ instead of the formally precise ‘a vertex of the link for $C$’. We define ‘the link for $D$’ as ‘the link for $C$’ with ‘$D$’ in place of ‘$C$’. Note that every edge in the separator of $(A,B)$ that is not diagonal lies in at most one of the edge-links for $C$ and $D$. We define the links for the sides $A$ and $B$ of $(A,B)$ analogously with the separations ‘$(A,B)$’ and ‘$(C,D)$’ interchanged. The link for a side $X$ is adjacent to the two corners for the pairs that contain the side $X$. Two links are adjacent if there is a corner they are both adjacent to. Every link is adjacent to all but one link; we refer to that link as its opposite link.
@@
Two mixed-separations $(A,B)$ and $(C,D)$ of a graph are nested if and only if they admit a corner such that it and the two adjacent links are empty. Thus, $(A,B)$ and $(C,D)$ cross as soon as two opposite links are nonempty.
@@
Indeed, we have $A\subseteq C$ and $B\supseteq D$ if and only if the corner for $\{A,D\}$ and its two adjacent links are empty. ∎
Section 1.3 offers an alternative definition of nestedness. We may and will use the two definitions interchangeably.
Suppose that we are given sides $X\in\{A,B\}$ and $Y\in\{C,D\}$. The corner-separator $L(X,Y)$ at the corner for $\{X,Y\}$ is the union of the two links adjacent to the corner for $\{X,Y\}$ together with the centre but without those diagonal edges that do not have an endvertex in the corner for $\{X,Y\}$; see Figure 1.3 for a picture. Two corner-separators are opposite or adjacent if their respective corners are.
Figure 1.3. $(A,B)$ and $(C,D)$ cross. The corner separator $L(A,C)$ contains all red vertices and edges.
An edge joining vertices in two opposite links is called a jumping edge; see Figure 1.4. It is straightforward to check that jumping edges are the only edges in the separators $S(A,B)$ and $S(C,D)$ that are not in any links or the centre and thus not in any corner separators.
@@
Let $a,b,c,d$ denote the sizes of the links for $A,B,C,D$, respectively. Let $x$ denote the size of the centre. Without loss of generality, the separators at the corners for $\{A,C\}$ and $\{B,D\}$ have size three. By Section 1.3, every diagonal edge has its endvertices in the corners for $\{A,C\}$ and $\{B,D\}$. Hence $a+c+x=3$ and $b+d+x=3$. Since $(A,B)$ and $(C,D)$ have order three, we further have $c+d+x=3$ and $a+b+x=3$. Considering the two equations that contain $a$, we find that $b=c$. Considering the two equations that contain $c$, we find that $a=d$. Without loss of generality, $a=d\leqslant b=c$.
@@
Suppose for a contradiction that this fails. Then two adjacent links have size one, and the other two links have size two, by Section 1.3. Hence the centre is empty. Let $X$ denote the corner whose adjacent links have size one. As the corner-separator for $X$ has size two, but $G$ is 3-connected, the corner $X$ is empty. Since $X$ is empty, not both adjacent links can consist of edges only, so one link contains a vertex $v$. Since the two links adjacent to $X$ have size one and the corner $X$ is empty, it follows that the vertex $v$ has at most one neighbour in one of the sides of the two crossing tri-separations. This contradicts the definition of tri-separation. ∎
@@
Without loss of generality, the separators at the corners for $\{A,C\}$ and $\{B,D\}$ have size three. Suppose for a contradiction that there is a diagonal edge $uv$. By Section 1.3, the ends $u$ and $v$ lie in the corners for $\{A,C\}$ and $\{B,D\}$, respectively say.
Claim 1.3.0.1.
@@
Proof of Claim. By Section 1.3, there are no jumping edges. Hence it suffices to show that all links have size $\ell=1$. Suppose for a contradiction that all links are empty, i.e. that $\ell=0$. Then the centre has size three. But since the centre contains the diagonal edge $uv$, the separators at the corners for $\{A,D\}$ and $\{B,C\}$ have size two. Then, since $G$ is 3-connected, the corners for $\{A,D\}$ and $\{B,C\}$ are empty. Since all links are empty as well, it follows that $(A,B)$ and $(C,D)$ are nested, contradicting our assumption that they cross. $\diamondsuit$
By Claim 1.3.0.1, all links and the centre have size one. So the centre only consists of the diagonal edge $uv$. Hence the separators at the two corners for $\{A,D\}$ and $\{B,C\}$ have size two. As $G$ is 3-connected, the two corners for $\{A,D\}$ and $\{B,C\}$ are empty. It follows that all four links contain no vertices, since any vertex in a link would fail to have two neighbours in some side of $(A,B)$ or $(C,D)$, contradicting that $(A,B)$ and $(C,D)$ are tri-separations. Hence all four links contain edges, and only edges. But then each of these edges must have an end in the corner for $\{A,D\}$ or $\{B,C\}$, contradicting that these corners are empty. ∎
Figure 1.4. Two tri-separations of a $K_{4}$ cross with two jumping edges (red and blue)
@@
This proof is supported by Figure 1.4. Since no two opposite corner-separators have size three, we find two adjacent corners whose separators have size at most two by Section 1.3. Say these are the corners $X$ and $Y$ for $\{A,C\}$ and $\{A,D\}$, respectively. As $G$ is 3-connected, both corners $X$ and $Y$ must be empty. Since $(A,B)$ is a tri-separation, $A\setminus B$ contains some vertex $a$. Since $X$ and $Y$ are empty, the vertex $a$ lies in the link for $A$. As $G$ is 3-connected, $a$ has degree at least three. Since $X$ and $Y$ are empty, and since the corner-separators at $X$ and $Y$ have size at most two, some edge incident to $a$ is a jumping edge. Let $b$ denote the other endvertex of this jumping edge, so $b$ lies in the link for $B$.
@@
A mixed-separation $(A,B)$ of a graph $G$ is half-connected if $G[A\setminus B]$ or $G[B\setminus A]$ are connected.
@@
In particular, there are no jumping edges. Moreover, if $(A,B)$ or $(C,D)$ is half-connected, then 1 holds.
@@
Since $G\neq K_{4}$, it follows from Section 1.3 that two opposite corner-separators have size three. Then all links have the same size $\ell\in\{0,1\}$ by Section 1.3. By Section 1.3, there are no jumping edges. By Section 1.3, there are no diagonal edges either. Hence the centre contains no edges, and the size of the centre is determined by $\ell$. If $\ell=0$, then the centre has size three; if $\ell=1$, then the centre has size one.
The ‘Moreover’ part follows from Section 1.3. ∎

### v3

@@ 3 Properties of tri-separations
The implication (2$\to$(1) is clear. For (1$\to$(2) suppose that $G[A]$ contains no cycle, say. We first show that the side $B$ cannot have exactly two vertices. Indeed, if $B$ has size two, then $G[B]$ has maximum degree at most one, and since $(A,B)$ is a tri-separation it follows that $A\cap B$ must be empty. Thus, $A$ and $B$ are the two sides of a cut of size three. Hence some vertex in $B$ has degree at most two in $G$, which contradicts 3-connectivity. So $B$ does not have size two. As we are done otherwise, from now on we assume that the side $B$ contains at least three vertices.
@@
Let any trivial tri-separation be given. By Section 3, it is of the form ($(\{v\},V\setminus\{v\})$, say. Then the vertex $v$ has degree three in $G$. Let $(C,D)$ be any strong tri-separation. Since $(C,D)$ is strong, the vertex $v$ is not in $C\cap D$. So it is in precisely one of $C\setminus D$ and $D\setminus C$, say in $C\setminus D$. Then $D\subseteq V\setminus\{v\}$, which gives ($(\{v\},V\setminus\{v\})\leqslant(C,D)$. ∎
@@
Let us show that no two edges in $S(A,B)$ share an end. For this, let us suppose for a contradiction that two edges $e,f\in S(A,B)$ share an endvertex $v\in A\setminus B$, say. Let $x$ be the remaining element of $S(A,B)$ besides $e,f$ if it is a vertex, and otherwise let $x$ denote the endvertex in $A\setminus B$ of the edge in $S(A,B)$ besides $e,f$. Let $O$ be a cycle in $G[A]$. Since $O$ has at least three vertices, one of them is distinct from $v$ and $x$, and so is not in $B\cup\{v,x\}$. Hence the pair $(A,B\cup\{v,x\})$ is a 2-separation of $G$ with separator equal to $\{v,x\}$. This contradicts the fact that $G$ is 3-connected. ∎
@@
Since $G$ is 3-connected, $v$ has neighbours in $X_{1}\setminus X_{2}$ and in $X_{2}\setminus X_{1}$, and $v$ has degree at least three. So if (1) fails, we have (2). ∎
@@
Let $(X_{1},X_{2})$ be a mixed 3-separation of a 3-connected graph. We obtain a tri-separation from $(X_{1},X_{2})$ by deleting vertices from $X_{1}$ or $X_{2}$, as follows. For every vertex $v\in X_{1}\cap X_{2}$ that has fewer than two neighbours in some side $X_{i}$, the index $i=:i(v)$ is unique, $v$ has a unique neighbour $x(v)$ in $X_{i}$ that lies in $X_{i}\setminus X_{3-i}$, and $v$ has two neighbours in $X_{3-i}\setminus X_{i}$ by Section 3. For both $j\in\{1,2\}$ we obtain $X_{j}^{\prime}$ from $X_{j}$ by deleting all vertices $v$ with $i(v)=j$. Then $(X_{1}^{\prime},X_{2}^{\prime})$ is a tri-separation of $G$. Every vertex $v\in X_{1}\cap X_{2}$ that was removed from some side is not in the separator of $(X_{1}^{\prime},X_{2}^{\prime})$, but instead the edge $\{v,x(v)\}$ is in $S(X_{1}^{\prime},X_{2}^{\prime})$. In this context, we say that $v$ was reduced to the edge $\{v,x(v)\}$. We call $(X_{1}^{\prime},X_{2}^{\prime})$ the reduction of $(X_{1},X_{2})$. Note that the reduction $(X_{1}^{\prime},X_{2}^{\prime})$ is nontrivial if $(X_{1},X_{2})$ is nontrivial.
Let $(A,B)$ be a mixed 3-separation of a graph $G$. A strengthening of $(A,B)$ is a mixed 3-separation $(A^{\prime},B^{\prime})$ that is obtained from $(A,B)$ by deleting at once all vertices of $A\cap B$ from one of the sides that have degree three in $G$ and have a neighbour in $A\cap B$, then taking a reduction. Note that every vertex deleted from $A\cap B$ is replaced by a unique edge in the separator, so the order of the separator stays the same.
@@
Let $(A,B)$ be a mixed 3-separation of a 3-connected graph $G$. For every edge $uv$ in $G$ with both ends $u,v$ in the separator of $(A,B)$, there is a strengthening $(A^{\prime},B^{\prime})$ of $(A,B)$ so that $u,v\in B^{\prime}$.
@@
Given $(A,B)$, we obtain $A^{\prime\prime}$ from $A$ by deleting all vertices that lie in $A\cap B$ and have degree three in $G$ and have some neighbour in $A\cap B$, and we put $B^{\prime\prime}:=B$. Then we let $(A^{\prime},B^{\prime})$ be the reduction of $(A^{\prime\prime},B^{\prime\prime})$. Suppose now that $uv$ is an edge with ends $u,v\in A\cap B$. Then $u$ and $v$ have two neighbours in $B=B^{\prime\prime}$, so $u,v\in B^{\prime}$. ∎
@@
(1$\rightarrow$(2). If $G$ is internally 4-connected or $G=K_{3,3}$, then every 3-separation $(A,B)$ of $G$ has a side ($A$, say) that induces a claw; in particular, $G[A]$ contains no cycle, so $(A,B)$ is trivial.
@@
Case 1: the induced subgraph $G[X]$ has no edges. Then $|A\setminus B|\geqslant 2$ and $|B\setminus A|\geqslant 2$. By Section 3, there is a strong tri-separation $(A^{\prime},B^{\prime})$ of $(A,B)$ with $A\setminus B\subseteq A^{\prime}$ and $B\setminus A\subseteq B^{\prime}$. Since $A\setminus B$ and $B\setminus A$ have size at least two, so have $A^{\prime}$ and $B^{\prime}$. Thus $(A^{\prime},B^{\prime})$ is nontrivial by Section 3.
Case 2: the induced subgraph $G[X]$ contains an edge $x_{1}x_{2}$. Recall that $|G|\geqslant 5$. The only 3-connected graphs on exactly five vertices that have a 3-separator are the 4-wheel and $K_{5}^{-}$ ($K_{5}$ minus one edge). The unique $3$-separator of $K_{5}^{-}$ is a triangle and thus it is the separator of a nontrivial tri-separation. The 4-wheel has two nontrivial strong tri-separations. So it remains to consider the case that $G$ has at least six vertices. By symmetry, we may assume that $|A\setminus B|\geqslant 2$. Since $x_{1}x_{2}$ is an edge of $G$ with both ends in $X=A\cap B$, we find a strong tri-separation $(A^{\prime},B^{\prime})$ of $G$ with $A\setminus B\subseteq A^{\prime}$ and $B\setminus A\subseteq B^{\prime}$ such that $x_{1},x_{2}\in B^{\prime}$. Hence, the side $B^{\prime}$ misses at most one vertex of $B$. As $|B|\geqslant 4$, this gives $|B^{\prime}|\geqslant 3$. We also have $|A^{\prime}|\geqslant|A\setminus B|$, and $|A\setminus B|\geqslant 2$ by assumption. Hence the strong tri-separation $(A^{\prime},B^{\prime})$ is nontrivial by Section 3. ∎
4 Nested or crossed: analysing corner diagrams
@@
Let $(A,B)$ and $(C,D)$ be mixed-separations of a graph $G$. The following definitions all depend on the context that $(A,B)$ and $(C,D)$ are given. They are supported by Figure 4, which is commonly referred to as a ‘corner diagram’.
Figure 4: $(A,B)$ and $(C,D)$ cross. Corners are blue, links are red, the centre is yellow.
@@
An edge $e$ in the separator of $(A,B)$ is in the edge-link for $C$ if it is not diagonal and has an endvertex in one of the corners for $C$; that is, in one of the corners for $\{A,C\}$ or $\{B,C\}$. The link for $C$ is the union of the edge-link for $C$ and the vertex set $(A\cap B)\setminus D$. In a slight abuse of notation we will sometimes say things like ‘a vertex of the link $(A\cap B)\setminus D$’ instead of the formally precise ‘a vertex of the link for $C$’. We define ‘the link for $D$’ as ‘the link for $C$’ with ‘$D$’ in place of ‘$C$’. Note that every edge in the separator of $(A,B)$ that is not diagonal lies in at most one of the edge-links for $C$ and $D$. We define the links for the sides $A$ and $B$ of $(A,B)$ analogously with the separations ‘$(A,B)$’ and ‘$(C,D)$’ interchanged. The link for a side $X$ is adjacent to the two corners for the pairs that contain the side $X$. Two links are adjacent if there is a corner they are both adjacent to. Every link is adjacent to all but one link; we refer to that link as its opposite link. Note that all four links are pairwise disjoint.
@@
Let $(A,B)$ and $(C,D)$ be two mixed-separations of a graph $G$. Assume that the corner for $\{A,C\}$ is empty. Let $L_{A}$ denote the link for $A$ and let $L_{C}$ denote the link for $C$. Then either there is a vertex in $L_{A}\cup L_{C}$ or $L_{A}=\emptyset=L_{C}$.
@@
Assume for a contradiction that $L_{A}\cup L_{C}\subseteq E(G)$ but $L_{A}$ is nonempty. Let $e\in L_{A}$ be an edge. By the definition of link, $e$ has an endvertex in the corner for $\{A,C\}$ or in $L_{C}$. The former is impossible as the corner for $\{A,C\}$ is empty by assumption. Hence $L_{C}$ contains an endvertex of $e$, a contradiction. ∎
[…]
Two mixed-separations $(A,B)$ and $(C,D)$ of a graph $G$ are nested if and only if they admit a corner such that it and the two adjacent links are empty. Thus, $(A,B)$ and $(C,D)$ cross as soon as two opposite links are nonempty.
Proof.
For the forward implication, suppose that $(A,B)\leqslant(C,D)$, say. Let $L_{A}$ denote the link for $A$, and let $V_{A}:=L_{A}\cap V(G)$. Similarly, let $L_{D}$ denote the link for $D$, and let $V_{D}:=L_{D}\cap V(G)$. Let $X$ denote the corner for $\{A,D\}$. From $A\subseteq C$ we get that $X$ and $V_{D}$ are empty. From $B\supseteq D$ we get that $X$ and $V_{A}$ are empty. Now Section 4 (with ‘$\{A,D\}$’ in place of ‘$\{A,C\}$’) gives $L_{A}=\emptyset=L_{D}$, as desired.
For the backward implication, suppose that $X=L_{A}=L_{D}=\emptyset$, say. In particular, $V_{A}$ and $V_{D}$ are empty. From $X$ and $V_{D}$ being empty we get $A\subseteq C$. From $X$ and $V_{A}$ being empty we get $B\supseteq D$. Hence $(A,B)\leqslant(C,D)$. ∎
Section 4 offers an alternative definition of nestedness. We may and will use the two definitions interchangeably.
Suppose that we are given sides $X\in\{A,B\}$ and $Y\in\{C,D\}$. The corner-separator $L(X,Y)$ at the corner for $\{X,Y\}$ is the union of the two links adjacent to the corner for $\{X,Y\}$ together with the centre but without those diagonal edges that do not have an endvertex in the corner for $\{X,Y\}$; see Figure 5 for a picture. Two corner-separators are opposite or adjacent if their respective corners are.
Figure 5: $(A,B)$ and $(C,D)$ cross. The corner separator $L(A,C)$ contains all red vertices and edges.
An edge joining vertices in two opposite links is called a jumping edge; see Figure 7. It is straightforward to check that jumping edges are the only edges in the separators $S(A,B)$ and $S(C,D)$ that are not in any links or the centre and thus not in any corner separators.
@@
Let $a,b,c,d$ denote the sizes of the links for $A,B,C,D$, respectively. Let $x$ denote the size of the centre. Without loss of generality, the separators at the corners for $\{A,C\}$ and $\{B,D\}$ have size three. By Section 4, every diagonal edge has its endvertices in the corners for $\{A,C\}$ and $\{B,D\}$. Hence $a+c+x=3$ and $b+d+x=3$. Since $(A,B)$ and $(C,D)$ have order three, we further have $c+d+x=3$ and $a+b+x=3$. Considering the two equations that contain $a$, we find that $b=c$. Considering the two equations that contain $c$, we find that $a=d$. Without loss of generality, $a=d\leqslant b=c$.
@@
Suppose for a contradiction that this fails. Then two adjacent links have size one, and the other two links have size two, by Section 4. See Figure 6. As any two opposite links have sizes one and two but are contained in a common separator of size three, the centre must be empty. Let $X$ denote the corner whose adjacent links have size one. Denote these two links by $L_{1}$ and $L_{2}$. As the corner-separator for $X$ has size at most two, but $G$ is 3-connected, the corner $X$ is empty. By Section 4, there is a vertex $v$ in $L_{1}$ or $L_{2}$. Since $L_{1}$ and $L_{2}$ have size one and the corner $X$ is empty, it follows that the vertex $v$ has at most one neighbour in one of the sides of the two crossing tri-separations. This contradicts the definition of tri-separation. ∎
Figure 6: The situation in the proof of Section 4
@@
Without loss of generality, the separators at the corners for $\{A,C\}$ and $\{B,D\}$ have size three. Suppose for a contradiction that there is a diagonal edge $uv$. By Section 4, the ends $u$ and $v$ lie in the corners for $\{A,C\}$ and $\{B,D\}$, respectively say.
Claim 4.0.1.
@@
Proof of Claim. By Section 4, there are no jumping edges. So the separator $S(A,B)$ is partitioned into two opposite links and the centre, and so is the separator $S(C,D)$. As both separators have size three, to show that all links and the centre have size one it suffices to show that all links have size $\ell=1$. Suppose for a contradiction that all links are empty, i.e. that $\ell=0$. Then the centre has size three. But since the centre contains the diagonal edge $uv$, the separators at the corners for $\{A,D\}$ and $\{B,C\}$ have size two. Then, since $G$ is 3-connected, the corners for $\{A,D\}$ and $\{B,C\}$ are empty (as otherwise the separator at such a corner is a mixed 2-separator of $G$ that separates a vertex in the corner from $u$ and $v$, contradicting 3-connectivity). Since all links are empty as well, it follows that $(A,B)$ and $(C,D)$ are nested, contradicting our assumption that they cross. $\diamondsuit$
By Claim 4.0.1, all links and the centre have size one. So the centre only consists of the diagonal edge $uv$. Hence the separators at the two corners for $\{A,D\}$ and $\{B,C\}$ have size two. As $G$ is 3-connected, the two corners for $\{A,D\}$ and $\{B,C\}$ are empty (as otherwise the separator at such a corner is a mixed 2-separator of $G$ that separates a vertex in the corner from $u$ and $v$). It follows that all four links contain no vertices, since any vertex in a link would fail to have two neighbours in some side of $(A,B)$ or $(C,D)$, contradicting that $(A,B)$ and $(C,D)$ are tri-separations. Hence all four links contain edges, and only edges. But then each of these edges must have an end in the corner for $\{A,D\}$ or $\{B,C\}$, contradicting that these corners are empty. ∎
Figure 7: Two tri-separations of a $K_{4}$ cross with two jumping edges (red and blue)
@@
This proof is supported by Figure 7. Since no two opposite corner-separators have size three, we find two adjacent corners whose separators have size at most two by Section 4. Say these are the corners $X$ and $Y$ for $\{A,C\}$ and $\{A,D\}$, respectively. As $G$ is 3-connected, both corners $X$ and $Y$ must be empty. Since $(A,B)$ is a tri-separation, it is also a mixed-separation, and by the definition of mixed-separation we have that $A\setminus B$ contains some vertex $a$. Since $X$ and $Y$ are empty, the vertex $a$ lies in the link for $A$. As $G$ is 3-connected, $a$ has degree at least three. Since $X$ and $Y$ are empty, and since the corner-separators at $X$ and $Y$ have size at most two, some edge incident to $a$ is a jumping edge. Let $b$ denote the other endvertex of this jumping edge, so $b$ lies in the link for $B$.
@@
A mixed-separation $(A,B)$ of a graph $G$ is half-connected if $G[A\setminus B]$ or $G[B\setminus A]$ is connected.
@@
In particular, there are no jumping edges. Moreover, if $(A,B)$ or $(C,D)$ is half-connected, then (1) holds.
@@
Since $G\neq K_{4}$, it follows from Section 4 that two opposite corner-separators have size three. Then all links have the same size $\ell\in\{0,1\}$ by Section 4. By Section 4, there are no jumping edges. By Section 4, there are no diagonal edges either. Hence the centre contains no edges, and the size of the centre is determined by $\ell$. If $\ell=0$, then the centre has size three; if $\ell=1$, then the centre has size one.
The ‘Moreover’ part follows from Section 4. ∎
[…]
Proof.
By Section 3, $(A,B)$ is nontrivial. Since $S(A,B)$ consists of edges, $(A,B)$ is strong. By Section 3, $(A,B)$ is nested with all trivial tri-separations of $G$. By the Nested or crossed: analysing corner diagrams (4), $(A,B)$ is nested with all nontrivial tri-separations of $G$. ∎